40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 21 09:02:03 2008, steve, Update, VAC, tp 2 failed 
    Reply  Fri Mar 21 11:54:38 2008, rob, Update, VAC, tp 2 failed 
Message ID: 388     Entry time: Fri Mar 21 09:02:03 2008     Reply to this: 389
Author: steve 
Type: Update 
Category: VAC 
Subject: tp 2 failed 
Small turbo #2 is the forepump of the maglev.
It failed last night, shut down the maglev and interlock closed V1
Ifo pressure is 20 mTorr now. The Yarm was still locked at 8am this morning.
The PSL beam to MC was blocked just before the output periscope.
The psl mechanial shutter did not work from epic screen.
ELOG V3.1.3-